# Role of quantum geometric tensor on wavepacket dynamics in two-dimensional non-Hermitian systems

### v1

@@ Abstract
Generalizing the QGT to systems effectively described by non-Hermitian Hamiltonians  [26, 27, 28, 29, 30, 31, 32, 33] is of fundamental importance for understanding the dynamical effects in systems with gain and loss. Non-Hermitian systems have attracted a great amount of interests because they can feature novel topological invariants [34], topological phases [34, 35], and topological edge states [36, 37, 38, 39, 40, 41, 42]. Non-Hermitian effects have also been experimentally studied in a variety of physical systems including cold atoms [43, 44], photonics [45, 46, 47, 48, 49], mechanical systems [50], electric circuits [51], magnons [52, 53], spin liquid [54], and exciton polaritons [55, 56].
@@
The second-order field-induced corrections to the anomalous non-Hermitian Berry connection has two components, i.e. $\operatorname{Re}[\mathbf{F}\cdot(\tilde{\mathbf{A}}_{00}^{RR}-\tilde{\mathbf{A}}_{00}^{LR})]=I_{0}+I_{1}+I_{2}$, where $I_{0}=\mathbf{F}\cdot(\mathbf{A}_{00}^{RR}-\mathbf{A}_{00}^{LR})]$ is the first-order term. The first correction is proportional to the difference of the RR and LR QMTs and has the form
@@
To demonstrate the relevance of the QGT components to the description of dynamical effects in a two-band non-Hermitian system, we simulate the wave-packet dynamics in an exciton-polariton system formed by strong coupling of an optically active, 2D-confined exciton to a photon confined in an optical microcavity. It is a 2D (two-dimensional) system featuring rich non-Hermitian effects due to inherent loss and gain [78]. Due to a pseudospin (polarisation) degree of freedom, it can also feature non-trivial bands with exceptional points in momentum space [56, 76, 73]. In principle, exciton-polaritons can provide an experimental platform for verifying the results presented here since components of both the QGT [9, 15] and its non-Hermitian generalization [73] can be measured in this system.
@@
We initialized the wave-packet at $\mathbf{k}=(0,0.9)\mu$m-1 in the upper eigenstate which has the larger $\operatorname{Im}E$ (or less loss) [see the dot in Figs. 1(c,d)]. The two bands are separated at this point in momentum space and $\partial_{k_{x}}\operatorname{Re}\epsilon_{0}=0$. An external force $\mathbf{F}=(0,5\times 10^{-4})$ eV/$\mu$m is applied accelerating the wave-packet along the $+k_{y}$-axis towards smaller (real) energy spacing. As the bands get closer, the mixing between the two eigenstates will be more profound and can result in oscillation due to the zitterbewegung effect [82, 75]. However, since the wave packet evolves into the less dissipative eigenstate, the inter-band mixing will gradually disappear and so will the oscillation. We also ensure that the wave packet does not cross the imaginary Fermi arc [56], where the imaginary parts of the eigenenergies cross, otherwise the single-band approximation will fail [76] as the loss rates of the eigenstates will switch. The numerical results are presented as dots in Figs. 2.
@@
The solution of Eq. (11) give a better fit to the numerical results than the solution of Eq. (6). This is apparent from the discrepancy between the numerical and analytical solutions plotted in Figs. 2 (c,d). Although the nonadiabatic correction is rather small in our example, a larger force or a smaller energy gap will result in a much larger correction, meaning that the previously derived equation of motion without perturbative corrections Eq. (6) will no longer be adequately describing the dynamics. In this case, however, we might also need to extend our treatment beyond the first-order perturbation theory and consider higher order corrections. On the other hand, if the wave packet propagates for a longer time, it would cross the imaginary Fermi arc, and the semi-classical equations of motion will no longer be valid [76].

### v2

# Quantum geometric tensor and wavepacket dynamics in two-dimensional non-Hermitian systems

@@ Abstract
Generalizing the QGT to systems effectively described by non-Hermitian Hamiltonians  [26, 27, 28, 29, 30, 31, 32, 33] is of fundamental importance for understanding the dynamical effects in systems with gain and loss. Non-Hermitian systems have attracted a great amount of interests because they can feature novel topological invariants [34], topological phases [34, 35], and topological edge states [36, 37, 38, 39, 40, 41, 42, 43, 44]. Non-Hermitian effects have also been experimentally studied in a variety of physical systems including cold atoms [45, 46], photonics [47, 48, 49, 50, 51], mechanical systems [52], electric circuits [53], magnons [54, 55], spin liquid [56], and exciton polaritons [57, 58].
@@
The second-order field-induced corrections to the anomalous non-Hermitian Berry connection have two components, i.e. $\operatorname{Re}[\mathbf{F}\cdot(\tilde{\mathbf{A}}_{00}^{RR}-\tilde{\mathbf{A}}_{00}^{LR})]=I_{0}+I_{1}+I_{2}$, where $I_{0}=\mathbf{F}\cdot(\mathbf{A}_{00}^{RR}-\mathbf{A}_{00}^{LR})]$ is the first-order term. The first correction is proportional to the difference of the RR and LR QMTs and has the form
@@
To demonstrate the relevance of the QGT components to the description of dynamical effects in a two-band non-Hermitian system, we simulate the wave-packet dynamics in an exciton-polariton system formed by strong coupling of an optically active, 2D-confined exciton to a photon confined in an optical microcavity. It is a 2D (two-dimensional) system featuring rich non-Hermitian effects due to inherent loss and gain [80]. The non-Hermiticity of the system arises from the decay of polaritons, mainly due to photons leaking out through the imperfect mirrors of the microcavity. Due to a pseudospin (polarisation) degree of freedom, this system can also feature non-trivial bands with exceptional points in momentum space [58, 78, 75]. In principle, exciton-polaritons can provide an experimental platform for verifying the results presented here since components of both the QGT [9, 15] and its non-Hermitian generalization [75] can be measured in this system.
@@
We initialized the wave-packet at $\mathbf{k}=(0,0.9)\mu$m-1 in the upper eigenstate which has the larger $\operatorname{Im}E$ (or less loss) [see the dot in Figs. 1(c,d)]. The two bands are separated at this point in momentum space and $\partial_{k_{x}}\operatorname{Re}\epsilon_{0}=0$. Experimentally, the wave-packet can be prepared using resonant optical excitation, where the exciting laser wavelength is adjusted to match the energy and the excitation angle is tuned to match the momentum of the target state [15].
An external force $\mathbf{F}=(0,5\times 10^{-4})$ eV/$\mu$m is applied accelerating the wave-packet along the $+k_{y}$-axis towards smaller (real) energy spacing. In an optical microcavity, this force naturally arises from the small variation or tilt in the cavity length introduced by the sample fabrication process, which translates to a linear energy gradient experienced by polaritons [85]. As the bands get closer, the mixing between the two eigenstates will be more profound and can result in oscillation due to the zitterbewegung effect [86, 77]. However, since the wave packet evolves into the less dissipative eigenstate, the inter-band mixing will gradually disappear and so will the oscillation. We also ensure that the wave packet does not cross the imaginary Fermi arc [58], where the imaginary parts of the eigenenergies cross, otherwise the single-band approximation will fail [78] as the loss rates of the eigenstates will switch. The numerical results are presented as dots in Figs. 2.
@@
The solution of Eq. (11) gives a better fit to the numerical results than the solution of Eq. (6). This is apparent from the discrepancy between the numerical and analytical solutions plotted in Figs. 2 (c,d). Although the nonadiabatic correction is rather small in our example, a larger force or a smaller energy gap will result in a much larger correction, meaning that the previously derived equation of motion without perturbative corrections Eq. (6) will no longer be adequately describing the dynamics. In this case, however, we might also need to extend our treatment beyond the first-order perturbation theory and consider higher order corrections. On the other hand, if the wave packet propagates for a longer time, it would cross the imaginary Fermi arc, and the semi-classical equations of motion will no longer be valid [78].